OpenStudy (anonymous):

How do you find f^-1 if f(x)=12x-10

OpenStudy (aravindg):

put y=12x-10 solve for x

OpenStudy (gorv):

f(x)=12x-10 f(x)+10=12x (f(x)+10)/12=x now for finding inverse interchange f(x) and x (x+10/12)= inverse of f(x)